OpenStudy (anonymous):

I forgot the operation, sqrt(4sqrt2)

OpenStudy (anonymous):

\[\sqrt{4\sqrt{2}}\]

OpenStudy (anonymous):

ok how do you find \[\sqrt{2\sqrt{2}}\]?

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

\[\sqrt{\sqrt{?}}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how do I find a sqrt of a sqrt?

OpenStudy (anonymous):